A ball is dropped from the top of the building having h=50m .from the 10m top of the building a window of length=2m is situated then find out the time period when ball crossed the window

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
Velocity after it descends by 10m till the start of the window = v
=>v2=u2+2gh
=>v2=2xgx10
=>v=10√2 m/s
 
S=ut+at2/2
=>2=10√2t+5t2/2
=>5t2+10√2t-2=0
=>t=(-10√2 +or- √(200+40) )/10 s
=>t = 0.13s
 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)